# cmaes : A Simple yet Practical Python Library for CMA-ES

### v2

@@ Abstract.
The covariance matrix adaptation evolution strategy (CMA-ES) has been highly effective in black-box continuous optimization, as demonstrated by its success in both benchmark problems and various real-world applications. To address the need for an accessible yet potent tool in this domain, we developed cmaes, a simple and practical Python library for CMA-ES. cmaes is characterized by its simplicity, offering intuitive use and high code readability. This makes it suitable for quickly using CMA-ES, as well as for educational purposes and seamless integration into other libraries. Despite its simplistic design, cmaes maintains enhanced functionality. It incorporates recent advancements in CMA-ES, such as learning rate adaptation for challenging scenarios, transfer learning, and mixed-integer optimization capabilities. These advanced features are accessible through a user-friendly API, ensuring that cmaes can be easily adopted in practical applications. We regard cmaes as the first choice for a Python CMA-ES library among practitioners. The software is available under the MIT license at https://github.com/CyberAgentAILab/cmaes.
@@
These authors contributed equally. This work was conducted while Masashi Shibata was affiliated with CyberAgent.
@@ 1. Introduction
Figure 1. Our library’s monthly download numbers have surpassed one million. Statistics are retrieved from PePy (Sincraian, 2018). Different colors represent different versions, with red indicating download numbers across all versions.
Black-box optimization focuses on optimizing an objective function whose internal structure is either unknown or inaccessible. In this approach, the objective function is regarded as a “black-box”: while it provides outputs in response to given inputs, the internal process generating these outputs remains concealed during optimization. This is invaluable when the objective function is complex, costly to evaluate, or lacks an explicit analytical form. The primary challenge in black-box optimization is enhancing the quality of the objective function value with minimal evaluations.
The covariance matrix adaptation evolution strategy (CMA-ES) (Hansen and Ostermeier, 2001; Hansen, 2016) stands out as a method for solving black-box continuous optimization problems. It optimizes by sampling candidate solutions from a multivariate Gaussian distribution, enabling parallel solution evaluations and showcasing significant parallelism. Its superiority is evident compared to other black-box optimization methods, particularly in challenging scenarios such as ill-conditioned, non-separable, or rugged problems (Rios and Sahinidis, 2013). The effectiveness of CMA-ES spans various real-world applications, including computer vision (Kikuchi et al., 2021b, a), natural language processing (Sun et al., 2022b, a), reinforcement learning (Ha and Schmidhuber, 2018), and automated machine learning (Nomura et al., 2021; Purucker and Beel, 2023).
In this work, we introduce cmaes, a Python-implemented CMA-ES library rapidly gaining popularity in the field. This is evidenced by over 250 GitHub stars and its integration into renowned Python libraries such as Optuna (Akiba et al., 2019) and Katib (George et al., 2020). Its popularity is further highlighted by approximately one million monthly downloads, as shown in Figure 1.
The cornerstone of cmaes lies in its simplicity and practicality: To ensure simplicity, our focus has been on high code readability, making it an inspiration for other evolution strategy libraries such as evojax (Tang et al., 2022) and evosax (Lange, 2023), and an excellent educational resource. Regarding its practicality, cmaes stands out by incorporating recent significant advancements in CMA-ES, setting it apart from other libraries. This integration includes (i) automatic learning rate adaptation for solving difficult problems such as multimodal and noisy environments without expensive hyperparameter tuning, (ii) transfer learning, and (iii) mixed-integer optimization. These advanced features are easily accessible through user-friendly APIs, establishing cmaes as the go-to choice for practitioners.
@@
In this paper, we aim to provide a comprehensive overview of cmaes’s motivation and distinctive features, making this paper a valuable guide for practitioners to effectively use the library. The remainder of this section discusses related work of Python software for CMA-ES. Section 2 explains the specifics of CMA-ES adopted in this study. Section 3 discusses the design philosophy behind cmaes, provides basic usage information, and discusses the software aspects of our library. Section 4 discusses several recent advances in the field of CMA-ES that can be used through cmaes. Section 5 concludes the paper with a summary and discussion.
@@
While there are several sophisticated libraries for CMA-ES, pycma (Hansen et al., 2019) is particularly renowned for its comprehensive features and detailed documentation. Indeed, pycma includes several features that our library lacks, such as handling of nonlinear constraints (Atamna et al., 2017), sophisticated restriction of the covariance matrix (Akimoto et al., 2014; Akimoto and Hansen, 2016, 2020), and surrogate-assisted methods (Hansen, 2019). However, it may pose challenges for users seeking an in-depth understanding due to its complexity. In contrast, our library cmaes focuses on basic and essential features, with an emphasis on simplicity and ease of understanding, appealing to practitioners who prefer straightforward implementations. This simplicity also enhances the library’s flexibility for integration of cmaes with other libraries.
Beyond pycma, the Python ecosystem houses other notable CMA-ES implementations. For example, evojax (Tang et al., 2022) and evosax (Lange, 2023) are JAX-based libraries, ideal for leveraging GPUs or TPUs in scalable optimization. pymoo (Blank and Deb, 2020) specializes in multi-objective optimization, an area not covered by our library. Additionally, Nevergrad (Rapin and Teytaud, 2018) offers a diverse range of optimization methods, including evolution strategies and Bayesian optimization, facilitating comparative studies between methods such as CMA-ES and Bayesian optimization (Frazier, 2018). This variety underscores the rich landscape of optimization tools available, each with its unique advantages and specialized applications.
@@ 2. CMA-ES
We consider minimization [^1] This is w.l.o.g. because replacing $f$ with $-f$ allows to consider maximization. of the objective function $f:\mathbb{R}^{d}\to\mathbb{R}$. CMA-ES optimizes $f$ by sampling solutions from a multivariate Gaussian distribution $\mathcal{N}(m,\sigma^{2}C)$, wherein $m\in\mathbb{R}^{d}$ is the mean vector, $\sigma\in\mathbb{R}_{>0}$ is the step-size, and $C\in\mathbb{R}^{d\times d}$ is the covariance matrix. Following the evaluation of solutions for $f$, CMA-ES updates the distribution parameters $m$, $\sigma$, and $C$ to produce more promising solutions. Figure 2 provides an example of optimization using CMA-ES. In this paper, we present a version of Ref. (Hansen, 2016) that our library cmaes adopts. For a detailed explanation of CMA-ES, see the Hansen’s excellent tutorial (Hansen, 2016).
CMA-ES begins by initializing the parameters $m^{(0)},\sigma^{(0)}$, and $C^{(0)}$ parameters. The following steps are then repeated until the stopping condition is met:
@@
In the $g+1^{\rm st}$ generation (with $g$ starting from 00), for the population size $\lambda$, $\lambda$ candidate solutions $x_{i}\ (i=1,2,\cdots,\lambda)$ are independently sampled from $\mathcal{N}(m^{(g)},(\sigma^{(g)})^{2}C^{(g)})$:
@@
(7) $\sigma^{(g+1)}=\sigma^{(g)}\exp\left(\frac{c_{\sigma}}{d_{\sigma}}\left(\frac{\|p_{\sigma}^{(g+1)}\|}{\mathbb{E}[\|\mathcal{N}(0,I)\|]}-1\right)\right),$
@@
(8) $\ +c_{1}\underbrace{\left[p_{c}^{(g+1)}\left(p_{c}^{(g+1)}\right)^{\top}-C^{(g)}\right]}_{\text{rank-one update}}+c_{\mu}\underbrace{\left[\sum_{i=1}^{\lambda}\left(w_{i}^{\circ}y_{i:\lambda}y_{i:\lambda}^{\top}-w_{i}C^{(g)}\right)\right]}_{\text{rank-$\mu$ update}},$
where $w_{i}^{\circ}:=w_{i}\cdot(1\text{ if }w_{i}\geq 0\text{ else }d/\|\sqrt{C^{(g)}}^{-1}y_{i:\lambda}\|^{2})$, $c_{m}$ is the learning rate for $m$; $c_{1}$ and $c_{\mu}$ are the learning rates for the rank-one and -$\mu$ updates of $C$, respectively; and $d_{\sigma}$ is the damping factor for the $\sigma$ adaptation. This adaptation of $\sigma$ is known as the cumulative step-size adaptation (CSA) (Hansen and Ostermeier, 2001). The update of distribution parameters in CMA-ES is closely related to the natural gradient descent (Akimoto et al., 2010).
@@
3. cmaes : Simplicity &\&& Practicality
This section aims to offer a comprehensive explanation of the various software aspects of our library. Initially, we delve into the core design philosophy of our library in Section 3.1, where we outline the fundamental principles that have shaped its development. Following this, Section 3.2 provides guidance on the installation process and the basic usage of the library, ensuring a smooth start for new users. We then shift our focus to features specifically designed to enhance the software’s quality: these include the continuous integration pipeline discussed in Section 3.3, the methodology for identifying unexpected errors via fuzzing in Section 3.4, and the animated visualization tools in Section 3.5. To demonstrate the practical application of cmaes in other libraries, Section 3.6 presents a use case involving Optuna (Akiba et al., 2019), showcasing the integration and utility of cmaes in a real-world context.
@@ 3.1. Design Philosophy
Our ultimate goal is to develop a Python CMA-ES library that becomes the first choice for practitioners. To achieve this, the library is designed with the dual aim of being both simple and practical, whose the detailed motivation is presented below.
@@
To achieve our goal, the library is designed with utmost simplicity. We identify three major benefits of this approach: Firstly, simplicity enhances software quality, easing debugging and maintenance, which is crucial for long-term reliability and practical use. Simplified maintenance also invites more feedback from the developer community, leading to a more refined product. Secondly, a simpler library is inherently more user-friendly. A straightforward API makes it easier for users to understand and utilize the library effectively. Lastly, from an educational standpoint, a simpler implementation aids in understanding the code in relation to the algorithm. This makes the library accessible not only to knowledgeable practitioners but also to beginners eager to learn about CMA-ES.
@@
While we prioritize keeping the library simple, it is essential to develop a library that solves the real-world challenges practitioners face. To maximize the library’s practicality, it is important to integrate insights from cutting-edge research on CMA-ES, particularly those with a strong practical motivation. However, it is equally important to ensure the library is user-friendly, as not all practitioners are knowledgeable about CMA-ES. The challenge we address is ensuring the practicality while maintaining the simplicity of our library.
@@
Based on our discussions, we have designed the library to be both simple and practical:
@@
This demonstrates the readability of our code and then its value for educational purposes. Our library has introduced animated visualizations (Section 3.5) that facilitates more an intuitive understanding of the behavior. In addition to the simplicity of the code itself, our APIs are also simple and easy-to-use, which will be presented in Section 3.2. To further be able to continuously keep the software quality, we have introduced continuous integration services (Section 3.3) and unexpected error identification testing via Fuzzing (Section 3.4).
To enhance practicality, our library has implemented three methods that are highly relevant in practical scenarios. The first method is for solving difficult problems such as multimodal and noisy issues without the need for expensive hyperparameter tuning. The second method is a transfer learning that accelerates CMA-ES by utilizing previous optimization results. The third method is for handling mixed-integer variables rather than only continuous variables. The details are explained in Section 4. To our knowledge, these methods have not yet been implemented into other evolution strategy libraries. Therefore, their presence in our library represents a distinctive feature. Importantly, these methods are generalizations of CMA-ES outlined in Section 2, and then do not require changes to the original CMA-ES implementation. Consequently, our library retains its simplicity. The API has also been designed for maximum simplicity, as will be detailed in Section 4.
@@ 3.2. Getting Started
This ask-and-tell interface, by decoupling the optimizer from the objective function, adds flexibility to the library. For example, it is easy to implement CMA-ES with restart strategies, such as IPOP-CMA-ES (Auger and Hansen, 2005) or BIPOP-CMA-ES (Hansen, 2009). Listing LABEL:list:ipop-cma-es provides an example code for IPOP-CMA-ES. In this example, the algorithm restarts with a new mean vector and a doubled population size when the stopping conditions are met (i.e., when should_stop() returns True). These stopping conditions are based on those used in pycma (Hansen et al., 2019). Additionally, we can specify box constraints by adding the argument, for example, bounds=np.array([[-3,3],[-3,3]]) to CMA($\cdot$); if a solution is not feasible, a resampling method is used, which is reasonable when the optimum solution is not very close to the constraint.
@@
3.3. Continuous Integration Pipeline
To maintain the quality of the library, we have utilized a continuous integration (CI) service. These services automatically activate in response to GitHub pull requests, facilitating software quality management without impeding the development pace. To achieve this, we have introduced several components into the development process: We perform tests across multiple Python versions to ensure compatibility. We employ static analysis tools for maintaining code quality. This includes Flake8 (Authority, 2010) for code style, Black (Foundation, 2018) for code formatting, and Mypy (Python, 2014) for type checking.
In addition to these basic validations, our repository conducts quick benchmarking, as illustrated in Figure 3. The aim of this quick benchmarking is to ensure that no significant bugs are introduced by recent changes. This benchmarking is to confirm whether the performance of cmaes is not quite different to that of pycma. For this, we use kurobako (Team, 2019) to create benchmark results efficiently. The benchmark results are uploaded onto Google Cloud Storage [^3] https://cloud.google.com/storage/ and subsequently incorporated into the corresponding pull request as comments, including images and descriptions. This integration is facilitated by custom GitHub Actions [^4] https://github.com/features/actions/.
It is important to acknowledge that the goal of our benchmarking process is to quickly identify evident bugs in the implementation without incurring substantial computational costs, rather than conducting rigorous benchmarking. For more comprehensive benchmarking, we recommend employing specialized software such as COCO (Hansen et al., 2021), a platform for comparing continuous optimizers in a black-box setting.
Figure 3. System diagram for CI pipeline.

### v3

@@ Abstract.
The covariance matrix adaptation evolution strategy (CMA-ES) has been highly effective in black-box continuous optimization, as demonstrated by its success in both benchmark problems and various real-world applications. To address the need for an accessible and powerful tool in this domain, we developed cmaes, a simple and practical Python library for CMA-ES. cmaes is characterized by its simplicity, offering intuitive use and high code readability. This makes it suitable for quick use of CMA-ES, as well as for educational purposes and seamless integration into other libraries. Despite its simple design, cmaes maintains advanced functionality. It incorporates recent advancements in CMA-ES, such as learning rate adaptation for challenging scenarios, transfer learning, mixed-variable optimization, and multi-objective optimization capabilities. These advanced features are accessible through a user-friendly API, ensuring that cmaes can be easily adopted in practical applications. We present cmaes as a strong candidate for a practical Python CMA-ES library aimed at practitioners. The software is available under the MIT license at https://github.com/CyberAgentAILab/cmaes.
@@ 1. Introduction
Black-box optimization focuses on optimizing an objective function whose internal structure is either unknown or inaccessible. In this approach, the objective function is regarded as a “black-box”: while it provides outputs in response to given inputs, the internal process generating these outputs remains concealed during optimization. This approach is invaluable when the objective function is complex, costly to evaluate, or lacks an explicit analytical form. The primary challenge in black-box optimization is enhancing the quality of the objective function value with minimal evaluations.
The covariance matrix adaptation evolution strategy (CMA-ES) (Hansen and Ostermeier, 2001; Hansen, 2016) stands out as a method for solving black-box continuous optimization problems. It optimizes by sampling candidate solutions from a multivariate Gaussian distribution, enabling parallel solution evaluations and showcasing significant parallelism. Its superiority is evident when compared to other black-box optimization methods, particularly in challenging scenarios such as ill-conditioned, non-separable, or rugged problems (Rios and Sahinidis, 2013). The effectiveness of CMA-ES spans various real-world applications, including computer vision (Kikuchi et al., 2021b, a), natural language processing (Sun et al., 2022b, a), reinforcement learning (Ha and Schmidhuber, 2018), model merging (Akiba et al., 2025), test-time adaptation (Niu et al., 2024), and automated machine learning (Nomura et al., 2021; Purucker and Beel, 2023).
In this work, we introduce cmaes, a Python-implemented CMA-ES library rapidly gaining popularity in the field. This is evidenced by over 450 GitHub stars and its integration into renowned Python libraries such as Optuna (Akiba et al., 2019) and Katib (George et al., 2020).
The cornerstone of cmaes lies in its simplicity and practicality: To ensure simplicity, our focus has been on high code readability, making it an inspiration for other evolution strategy libraries such as evojax (Tang et al., 2022) and evosax (Lange, 2023), a useful educational resource. Regarding its practicality, cmaes stands out by incorporating recent significant advancements in CMA-ES, setting it apart from other libraries. This integration includes (i) automatic learning rate adaptation for solving difficult problems such as multimodal and noisy environments without expensive hyperparameter tuning, (ii) transfer learning, (iii) mixed-variable optimization, and (iv) multi-objective optimization. These advanced features are easily accessible through user-friendly APIs, establishing cmaes as the primary choice for practitioners.
@@
In this paper, we aim to provide a comprehensive overview of cmaes’s motivation and distinctive features, making this paper a valuable guide for practitioners to effectively use the library. The remainder of this section discusses related work on Python software for CMA-ES. Section 2 explains the specifics of CMA-ES adopted in this study. Section 3 discusses the design philosophy behind cmaes, provides basic usage information, and discusses the software aspects of our library. Section 4 discusses several recent advances in the field of CMA-ES that can be used through cmaes. Section 5 concludes the paper with a summary and discussion.
@@
While there are several sophisticated libraries for CMA-ES, pycma (Hansen et al., 2019) is particularly renowned for its comprehensive features and detailed documentation. Indeed, pycma includes several features that our library lacks, such as handling of nonlinear constraints (Atamna et al., 2017), sophisticated restriction of the covariance matrix (Akimoto et al., 2014; Akimoto and Hansen, 2016, 2020), and surrogate-assisted methods (Hansen, 2019). However, it might pose challenges for users seeking an in-depth understanding due to its complexity. In contrast, our library cmaes focuses on basic and essential features, with an emphasis on simplicity and ease of understanding, appealing to practitioners who prefer straightforward implementations. This simplicity also enhances the library’s flexibility for integration of cmaes with other libraries.
Beyond pycma, the Python ecosystem houses other notable CMA-ES implementations. For example, evojax (Tang et al., 2022) and evosax (Lange, 2023) are JAX-based libraries, ideal for leveraging GPUs or TPUs in scalable optimization. pymoo (Blank and Deb, 2020) specializes in multi-objective optimization, which is only partially covered by our library. Additionally, Nevergrad (Rapin and Teytaud, 2018) offers a diverse range of optimization methods, facilitating comparative studies between methods such as CMA-ES and Bayesian optimization (Frazier, 2018). This variety underscores the rich landscape of optimization tools available, each with its unique advantages and specialized applications.
@@ 2. CMA-ES
We consider minimization [^1] This is w.l.o.g. because replacing $f$ with f-f allows to consider maximization. of the objective function $f:\mathbb{R}^{d}\to\mathbb{R}$. CMA-ES optimizes $f$ by sampling solutions from a multivariate Gaussian distribution $\mathcal{N}(m,\sigma^{2}C)$, wherein $m\in\mathbb{R}^{d}$ is the mean vector, $\sigma\in\mathbb{R}_{>0}$ is the step-size, and $C\in\mathbb{R}^{d\times d}$ is the covariance matrix. Following the evaluation of solutions for $f$, CMA-ES updates the distribution parameters $m$, $\sigma$, and $C$ to produce more promising solutions. Figure 1 provides an example of optimization using CMA-ES. In this paper, we present a version of Ref. (Hansen, 2016) that our library cmaes adopts. For a detailed explanation of CMA-ES, see the Hansen’s excellent tutorial (Hansen, 2016).
We briefly summarize CMA-ES to clarify the exact variant implemented in cmaes. CMA-ES begins by initializing the parameters $m^{(0)},\sigma^{(0)}$, and $C^{(0)}$ parameters. The following steps are then repeated until the stopping condition is met:
@@
In the $g+1^{\rm st}$ generation (with $g$ starting from 0), for the population size $\lambda$, $\lambda$ candidate solutions $x_{i}\ (i=1,2,\cdots,\lambda)$ are independently sampled from $\mathcal{N}(m^{(g)},(\sigma^{(g)})^{2}C^{(g)})$:
@@
(7) $\sigma^{(g+1)}=\sigma^{(g)}\exp\left(1,\frac{c_{\sigma}}{d_{\sigma}}\left(\frac{\|p_{\sigma}^{(g+1)}\|}{\mathbb{E}[\|\mathcal{N}(0,I)\|]}-1\right)\right),$
@@
(8) $\ +c_{1}\underbrace{\left[p_{c}^{(g+1)}\left(p_{c}^{(g+1)}\right)^{\top}-C^{(g)}\right]}_{\text{rank-one update}}+c_{\mu}\underbrace{\sum_{i=1}^{\lambda}w_{i}^{\circ}\left[y_{i:\lambda}y_{i:\lambda}^{\top}-C^{(g)}\right]}_{\text{rank-$\mu$ update}},$
where $w_{i}^{\circ}:=w_{i}\cdot(1\text{ if }w_{i}\geq 0\text{ else }d/\|\sqrt{C^{(g)}}^{-1}y_{i:\lambda}\|^{2})$, $c_{m}$ is the learning rate for $m$; $c_{1}$ and $c_{\mu}$ are the learning rates for the rank-one and -$\mu$ updates of $C$, respectively; and $d_{\sigma}$ is the damping factor for the $\sigma$ adaptation. The update of distribution parameters in CMA-ES is closely related to the natural gradient descent (Akimoto et al., 2010).
@@
3. cmaes: Simplicity &\& Practicality
This section aims to offer a comprehensive explanation of the various software aspects of our library. Initially, we delve into the core design philosophy of our library in Section 3.1, where we outline the fundamental principles that have shaped its development. Following this, Section 3.2 provides guidance on the installation process and the basic usage of the library, ensuring a smooth start for new users. We then shift our focus to features specifically designed to enhance the software’s quality: these include the methodology for identifying unexpected errors via fuzzing in Section 3.3, and the animated visualization tools in Section 3.4. To demonstrate the practical application of cmaes in other libraries, Section 3.5 presents a use case involving Optuna (Akiba et al., 2019), showcasing the integration and utility of cmaes in a real-world context.
@@ 3.1. Design Philosophy
Our ultimate goal is to develop a Python CMA-ES library that serves as a strong practical choice for real-world users. To this end, the library is designed with simplicity and practicality as its guiding principles.
@@
To achieve our goal, the library is designed with utmost simplicity. We identify three major benefits of this approach: Firstly, simplicity enhances software quality, easing debugging and maintenance, which is crucial for long-term reliability and practical use. Simplified maintenance also invites more feedback from the developer community, leading to a more refined product. Secondly, a more lightweight library is inherently more user-friendly. A straightforward API makes it easier for users to understand and utilize the library effectively. Lastly, from an educational standpoint, a simpler implementation aids in understanding the code in relation to the algorithm. This makes the library accessible not only to knowledgeable practitioners but also to beginners eager to learn about CMA-ES.
@@
While we prioritize keeping the library simple, it is essential to develop a library that solves the real-world challenges practitioners face. To maximize the library’s applicability in real-world settings, it is important to integrate insights from cutting-edge research on CMA-ES, particularly those with a strong practical motivation. However, it is equally important to ensure the library is user-friendly, as not all practitioners are knowledgeable about CMA-ES. The challenge we address is ensuring the practicality while maintaining the simplicity of our library.
@@
Based on the above principles, we have designed the library as follows:
@@
This demonstrates the readability of our code and its value for educational purposes. Our library has introduced animated visualizations (Section 3.4) that facilitate a more intuitive understanding of the behavior. In addition to the simplicity of the code itself, our APIs are also simple and easy-to-use, which will be presented in Section 3.2. To continuously maintain the software quality, we have implemented fuzz testing to identify unexpected errors (Section 3.3).
To enhance practicality, our library implements four methods that are highly relevant in practical scenarios. The first method is for solving difficult problems such as multimodal and noisy issues without the need for expensive hyperparameter tuning. The second method is a transfer learning that accelerates CMA-ES by utilizing previous optimization results. The third method is for handling mixed-variable domains, including integer and categorical variables rather than only continuous variables. The fourth method is a multi-objective optimizer that supports mixed-variable domains. The details are explained in Section 4. To the best of our knowledge, these methods have not yet been implemented in other evolution strategy libraries. Therefore, their presence in our library represents a distinctive feature. Importantly, these methods are generalizations of CMA-ES outlined in Section 2, and thus do not require changes to the original CMA-ES implementation. Consequently, our library retains its simplicity. The API has also been designed for maximum simplicity, as will be detailed in Section 4.
@@ 3.2. Getting Started
This ask-and-tell interface, by decoupling the optimizer from the objective function, adds flexibility to the library. For example, it is easy to implement CMA-ES with restart strategies, such as IPOP-CMA-ES (Auger and Hansen, 2005) or BIPOP-CMA-ES (Hansen, 2009). Listing LABEL:list:ipop-cma-es provides an example code for IPOP-CMA-ES. In this example, the algorithm restarts with a new mean vector and a doubled population size when the stopping conditions are met (i.e., when should_stop() returns True). These stopping conditions are based on those used in pycma (Hansen et al., 2019). Additionally, we can specify box constraints by adding the argument, for example, bounds=np.array([[-3,3],[-3,3]]) to CMA($\cdot$); if a solution is not feasible, a resampling method is used, which is reasonable when the optimum solution is not very close to the constraints.